# Autoregressive Models in Vision: A Survey

### v1

@@ Abstract
Autoregressive modeling has been a huge success in the field of natural language processing (NLP). Recently, autoregressive models have emerged as a significant area of focus in computer vision, where they excel in producing high-quality visual content. Autoregressive models in NLP typically operate on subword tokens. However, the representation strategy in computer vision can vary in different levels, i.e., pixel-level, token-level, or scale-level, reflecting the diverse and hierarchical nature of visual data compared to the sequential structure of language. This survey comprehensively examines the literature on autoregressive models applied to vision. To improve readability for researchers from diverse research backgrounds, we start with preliminary sequence representation and modeling in vision. Next, we divide the fundamental frameworks of visual autoregressive models into three general sub-categories, including pixel-based, token-based, and scale-based models based on the strategy of representation. We then explore the interconnections between autoregressive models and other generative models. Furthermore, we present a multi-faceted categorization of autoregressive models in computer vision, including image generation, video generation, 3D generation, and multi-modal generation. We also elaborate on their applications in diverse domains, including emerging domains such as embodied AI and 3D medical AI, with about 250 related references. Finally, we highlight the current challenges to autoregressive models in vision with suggestions about potential research directions. We have also set up a Github repository to organize the papers included in this survey at: https://github.com/ChaofanTao/Autoregressive-Models-in-Vision-Survey. [^1] Literature reviewed is current as of October 31, 2024, with updates planned.
@@ 1 Introduction
The timeline in Figure 1 illustrates the key milestones and developments in the evolution of visual autoregressive models, highlighting their transition from NLP to computer vision. To date, autoregressive models have been applied to a wide array of generative tasks, including image generation (Parmar et al., 2018; Chen et al., 2020), image super-resolution (Guo et al., 2022; Li et al., 2016), image editing (Yao et al., 2022; Crowson et al., 2022), image-to-image translation (Li et al., 2024e; d) and video generation (Tulyakov et al., 2018; Hong et al., 2022), multi-modal tasks (Yu et al., 2023c; Lu et al., 2022) and medical tasks (Ren et al., 2024; Tudosiu et al., 2024), .etc. This broad applicability underscores the potential for further exploration and application of autoregressive models.
@@
We emphasize that there are at least three distinct categories of visual autoregressive models defined by their sequence representation strategies: pixel-based, token-based, and scale-based models. Pixel-RNN  (Van Den Oord et al., 2016), as a representative pixel-wise model in pioneering next-pixel prediction by transforming a 2D image into a 1D pixel sequence, capturing both local and long-range dependencies but with high computational cost. Next-token prediction, inspired by NLP, compresses images into discrete tokens for efficient high-resolution processing, exemplified by models like VQ-VAE (Van Den Oord et al., 2017). VAR (Tian et al., 2024) introduces next-scale prediction, a hierarchical method generating content across multiple scales from coarse to fine autoregressively, capturing visual information at multiple resolutions. Each category offers unique advantages and challenges, making them promising directions for future research.
We further introduce a multi-perspective categorization of autoregressive models applied to computer vision, classifying existing models based on criteria such as the sequence representation strategy, the underlying framework, or the target task. Our categorization aims to provide a structured overview of how these models are utilized across various vision tasks. We then present both quantitative and qualitative metrics to assess their performance and applicability. Finally, we highlight the current limitations of autoregressive models, such as computational complexity and mode collapse, and propose potential directions for future research. In summary, this survey makes several contributions:
@@
RQ-VAE (Lee et al., 2022a), MoVQ (Zheng et al., 2022), DQ (Huang et al., 2023)
@@
DCT-Transformer (Nash et al., 2021), MAGVIT (Yu et al., 2023a),
MAGVIT-v2 (Yu et al., 2023b), Open-MAGVIT2 (Luo et al., 2024),
OmniTokenizer (Wang et al., 2024b), , fill=orange!10, text width=34em, align=left] [VQ-VAE (Van Den Oord et al., 2017), VQ-VAE-2 (Razavi et al., 2019),
@@
AiM (Li et al., 2024b), ZigMa Hu et al. (2024), DiM (Teng et al., 2024),
DiG (Zhu et al., 2024c), Diffusion-RWKV (Fei et al., 2024), fill=orange!10, text width=34em, align=left] ] [Scale-wise Generation3.1.1.14), fill=orange!10 , text width=20em [VAR (Zhang et al., 2024a) , fill=orange!10, text width=34em] ] ] [Text-to-Image Synthesis3.1.2), fill=green!1 , text width=20em [Token-wise Generation6), fill=green!1, text width=20em [DALL·E (Ramesh et al., 2021), CogView (Ding et al., 2021),
@@
MVG (Ren et al., 2024), fill=yellow!10, text width=34em] ] ] ] [Video Generation3.2), fill=pink!5 , text width=15em [Unconditional Video Generation3.2.1), fill=pink!5, text width=20em [VPNs (Kalchbrenner et al., 2017), MoCoGAN (Tulyakov et al., 2018),
@@
ST-LLM (Liu et al., 2025), Pyramid Flow (Jin et al., 2024a), fill=pink!5, text width=34em, align=left] [MAGE (Hu et al., 2022), VideoPoet (Kondratyuk et al., 2024) , fill=pink!5, text width=34em] ] [Embodied AI3.2.3), fill=pink!5 ,text width=20em [IRIS (Micheli et al., 2023), iVideoGPT (Wu et al., 2024d),
@@
ImAM (Qian et al., 2024), Argus3D (Qian et al., 2024) , fill=violet!10, text width=34em ] ] [Scene Generation3.3.3), fill=violet!10, text width=20em [Make-A-Scene (Gafni et al., 2022), SceneScript (Avetisyan et al., 2024) , fill=violet!10, text width=34em ] ] [3D Medical Generation3.3.4), fill=violet!10, text width=20em [SynthAnatomy (Tudosiu et al., 2022), BrainSynth (Tudosiu et al., 2024),
@@
Unalign (Corona-Figueroa et al., 2023), AutoSeq (Wang et al., 2024c), fill=violet!10, text width=34em, align=left ] ] ] [Multimodal Generation3.4), fill=teal!7, text width=18em [Understanding Framework3.4.1), fill=teal!7, text width=20em [BEiT  (Bao et al., 2021), BEiT-v2 (Peng et al., 2022), LLaVA  (Liu et al., 2024d),
VL-BEiT (Bao et al., 2022), BEiT-v3 (Wang et al., 2022b) , fill=teal!7, text width=34em, align=left ] ] [Unified Framework3.4.2), fill=teal!7, text width=20em [OFA (Wang et al., 2022a), CogView (Ding et al., 2021), M6 (Lin et al., 2021),
@@
Video-LaViT (Jin et al., 2024b), X-ViLA (Ye et al., 2024), fill=teal!7, text width=34em, align=left] [Chameleon (Team, 2024), Transfusion (Zhou et al., 2024)
@@ 2.1 Preliminary
The visual data is first transformed into a sequence of discrete elements. These elements may correspond to pixels, image patches, or latent codes derived from the visual content, depending on the specific model architecture. This transformation allows the visual data to be framed as a sequential modeling problem, akin to text generation in natural language processing.
@@ 2.2 Generic Frameworks
In the preceding sections, we have provided a general overview of visual autoregressive models about the essential concept of sequence representation and sequence modeling. Then, we consider the representation strategy to classify the visual autoregressive models further, i.e. pixel-based models, token-based models, scale-based models. Since representation strategy directly reflects how models handle the underlying structure and features of visual data. Different representation strategies influence the model’s granularity, efficiency, and applicability. Pixel-based models (Sec. 2.2.1) generate images pixel by pixel, capturing the most fine-grained spatial details. Token-based models (Sec. 2.2.2) represent images as a set of high-level discrete visual tokens, analogous to words in NLP. Scale-based models (Sec. 2.2.3) generate images through a multi-scale representation, allowing the model to handle visual content from different resolutions. We detail these three types of generic framework below.
@@ 2.2.2 Token-based Models
This quantization process enables a compact and discrete representation of the latent feature space, which is particularly advantageous in reducing the computational burden for high-dimensional image generation tasks. A seminal work on discrete representation is VQVAE (Van Den Oord et al., 2017), which introduced a two-stage paradigm that has become the de-facto standard for autoregressive vision generation. In this paradigm, an encoder-decoder architecture is initially trained to learn the discrete image representation. The encoder $E$ maps the input image $x$ to a latent space $z_{e}(x)$, where the continuous latent features are quantized into discrete codes using VQ. The decoder $D$ then reconstructs the image from these discrete codes. The training objective consists of a reconstruction loss and a commitment loss to ensure the latent codes effectively represent the input:
@@ 2.3 Relation to Other Generative Models
Variational Autoencoders a(VAEs) (Kingma, 2013) are a class of generative models that learn to map data into a continuous, lower-dimensional latent space and subsequently reconstruct it back to the original data space. This process is governed by optimizing a variational lower bound of the data likelihood. In contrast, autoregressive models directly capture the full data distribution by predicting each element sequentially, optimizing negative log-likelihood(NLL) as the training objective. While VAEs provide efficient sampling, their reliance on variational approximations often results in less sharp outputs (Child, 2021) and may suffer from posterior collapse (Zhao et al., 2019; Chen et al., 2017) when equipped with a powerful decoder. Autoregressive models, on the other hand, generate high-quality samples by modeling data dependencies at the original dimensionality, but they are slower in inference due to their sequential generation process. To address these limitations, hybrid approaches (Gulrajani et al., 2017; Van Den Oord et al., 2017; Chen et al., 2017) have been proposed that integrate the autoregressive process into VAEs. One prominent example is VQ-VAE (Van Den Oord et al., 2017), which utilizes a VAE framework to learn discrete latent spaces and autoregressive models to refine generation, effectively leveraging the strengths of both models for efficient and high-quality image synthesis.
@@
GANs (Goodfellow et al., 2014) are known for their ability to generate high-quality images with fast inference, enabled by their one-shot generation process. GANs demonstrate particular strength in domains with structured data, such as facial synthesis, where their low-dimensional latent space can be effectively leveraged for manipulating visual attributes (Xia et al., 2022). However, the adversarial training objective that GANs rely on often results in training instability and mode collapse (Liu et al., 2020), requiring a delicate balance between the generator and discriminator. Additionally, the adversarial training paradigm that underpins GANs can hinder their scalability to more diverse datasets and larger model sizes. In contrast, autoregressive models employ likelihood-based training, which ensures a stable training process. Despite their relatively slow sampling speeds, autoregressive models exhibit favorable scaling laws, where model performance consistently improves with larger datasets and increased model sizes (Kaplan et al., 2020). These properties make autoregressive models particularly well-suited for building universal generative models, and can be adapted to a wide range of downstream tasks throught zero-shot generalization of supervised fine-tuning(SFT), enabling flexibility and robustness across different applications.
@@
Diffusion models (Ho et al., 2020) have emerged as the state-of-the-art generation paradigm for a wide range of vision generation tasks. As recent challengers, autoregressive models share several characteristics with diffusion models. Both methods are capable of generating diverse, high-quality samples, yet both also suffer from inefficient inference due to their iterative or sequential generation processes. Additionally, both approaches are likelihood-based and optimize objectives such as Negative Log-Likelihood (NLL) or Evidence Lower Bound (ELBO), making them relatively easy to train. Recent advances (Esser et al., 2021b; Rombach et al., 2022) in both fields have focused on compressing visual content into latent spaces to improve the efficiency of high-resolution generation. Despite these similarities, there are fundamental differences in their generative paradigms. Diffusion models employ a predefined forward process that gradually corrupts data and an iterative denoising process to recover samples. This iterative process retains spatial locality at each step, which benefits visual coherence. However, the necessity to corrupt the training data with Gaussian noise may potentially limit their scalability for understanding tasks. Additionally, while diffusion models operate primarily in continuous spaces, their discrete variants still lag behind in performance, posing challenges for multimodal applications like text generation. Autoregressive models, by contrast, inherently introduce unidirectional biases and discretization, which might not be ideal for visual tasks, partially explaining their lag behind diffusion models in recent benchmarks. However, autoregressive models offer flexibility in handling diverse modalities and combining generation with understanding. This adaptability aligns with the emerging trend of integrating autoregressive models with Large Language Models (LLMs) to create a unified framework for multimodal input-output tasks, bridging both generation and understanding capabilities. These distinctions and complementarities between the two approaches suggest potential avenues for future research. Recent efforts have aimed to combine the strengths of both approaches, such as integrating bidirectional contexts (Esser et al., 2021a; Zhou et al., 2024; Xie et al., 2024b), autoregressive models with continuous representations (Li et al., 2024c; Zhou et al., 2024), discrete diffusion models (Gu et al., 2022), and using autoregressive models to enhance the understanding capabilities of diffusion models (Gu et al., 2024).
@@
3 Visual Autoregressive Models
3.1 Image Generation
3.1.1 Unconditional Image Generation
Pixel-wise Generation In pixel-wise generation, the model constructs an image pixel on a pixel-by-pixel basis, unconditionally predicting each pixel’s quantified value based solely on the sequence of previously generated pixels. This approach allows for the direct optimization of the likelihood function, thus producing highly detailed outputs, albeit at a significant computational cost. This unconditional generation marks the inception of AR generation research, laying the foundational paradigm upon which all subsequent work in this domain has been built.
3.1.1.1 Recurrent and Parallel Pixel Generation Techniques.
PixelRNN (Van Den Oord et al., 2016) uses RNNs to predict pixel values sequentially, effectively capturing complex dependencies between pixels. This work is the first to propose the use of an autoregressive approach for image generation. Building upon its success, PixelCNN (Van den Oord et al., 2016) adopts dilated convolutions to efficiently capture long-range pixel dependencies in image generation. PixelCNN++ (Salimans et al., 2017) optimizes PixelCNN further by using a discretized logistic mixture likelihood function and architectural improvements, enhancing image quality and efficiency. Reed et al. (2017) introduces a parallelized version of PixelCNN, reducing computational complexity and enabling faster high-resolution image generation.
3.1.1.2 Transformer-based Approaches for Pixel Generation.
Traditional CNNs and RNNs face scaling issues when used for autoregressive image generation on large datasets, primarily due to their limited ability to model long-term dependencies and their model architectures also struggle to scale up in size. To address this issue, Transformer (Vaswani, 2017) has been adapted for pixel-wise generation, leveraging its superior capacity to capture long-range dependencies within the data. PixelSNAIL (Chen et al., 2018) is the first work to combine causal convolutions with self-attention mechanisms, allowing the model to consider both local and broader contexts while preserving the sequential nature of pixel generation. Inspired by it, The Image Transformer (Parmar et al., 2018) employs self-attention to handle local neighborhoods in larger images, improving performance on complex tasks. ImageGPT (Chen et al., 2020) utilizes a decoder-only architecture and goes further by treating images as sequences of pixels, employing an autoregressive framework similar to that used in natural language processing. The above approaches demonstrate the power of Transformers in modeling long-range dependencies between pixels, laying the groundwork for future advancements in large-scale unsupervised image learning.
Token-wise Generation
Token-wise generation utilizes a different approach, where the image is divided into a sequence of tokens, and each token is generated based on the previously generated tokens, similar to how text is processed in natural language models. This paradigm often leverages models like transformers, which can handle larger contexts in one step compared to pixel-by-pixel generation. These tokens typically represent local patches of the image in a more compact form, allowing for faster generation and scalability in terms of image size and complexity. Token-wise generation focuses on the representation of images as sequences of discrete tokens, i.e., the design of image tokenizer, as well as the modeling of autoregressive generation process at token level.
3.1.1.3 (1) Design of Image Tokenizer

### v2

@@ Abstract
Autoregressive modeling has been a huge success in the field of natural language processing (NLP). Recently, autoregressive models have emerged as a significant area of focus in computer vision, where they excel in producing high-quality visual content. Autoregressive models in NLP typically operate on subword tokens. However, the representation strategy in computer vision can vary in different levels, i.e., pixel-level, token-level, or scale-level, reflecting the diverse and hierarchical nature of visual data compared to the sequential structure of language. This survey comprehensively examines the literature on autoregressive models applied to vision. To improve readability for researchers from diverse research backgrounds, we start with preliminary sequence representation and modeling in vision. Next, we divide the fundamental frameworks of visual autoregressive models into three general sub-categories, including pixel-based, token-based, and scale-based models based on the representation strategy. We then explore the interconnections between autoregressive models and other generative models. Furthermore, we present a multifaceted categorization of autoregressive models in computer vision, including image generation, video generation, 3D generation, and multimodal generation. We also elaborate on their applications in diverse domains, including emerging domains such as embodied AI and 3D medical AI, with about 250 related references. Finally, we highlight the current challenges to autoregressive models in vision with suggestions about potential research directions. We have also set up a Github repository to organize the papers included in this survey at: https://github.com/ChaofanTao/Autoregressive-Models-in-Vision-Survey.
@@ 1 Introduction
The timeline in Figure 1 illustrates the key milestones and developments in the evolution of visual autoregressive models, highlighting their transition from NLP to computer vision. To date, autoregressive models have been applied to a wide array of generative tasks, including image generation (Parmar et al., 2018; Chen et al., 2020), image super-resolution (Guo et al., 2022; Li et al., 2016), image editing (Yao et al., 2022; Crowson et al., 2022), image-to-image translation (Li et al., 2024e; d) and video generation (Tulyakov et al., 2018; Hong et al., 2022), multi-modal tasks (Yu et al., 2023c; Lu et al., 2022) and medical tasks (Ren et al., 2024a; Tudosiu et al., 2024), .etc. This broad applicability underscores the potential for further exploration and application of autoregressive models.
@@
We emphasize that there are at least three distinct categories of visual autoregressive models defined by their sequence representation strategies: pixel-based, token-based, and scale-based models. Pixel-RNN  (Van Den Oord et al., 2016), as a representative pixel-wise model in the pioneering of next-pixel prediction by transforming a 2D image into a 1D pixel sequence, capturing both local and long-range dependencies but with high computational cost. Next-token prediction, inspired by NLP, compresses images into discrete tokens for efficient high-resolution processing, exemplified by models like VQ-VAE (Van Den Oord et al., 2017). VAR (Tian et al., 2024) introduces next-scale prediction, a hierarchical method that generates content across multiple scales, from coarse to fine autoregressively capturing visual information at multiple resolutions. Each category offers unique advantages and challenges, making them promising directions for future research.
We further introduce a multi-perspective categorization of autoregressive models applied to computer vision, which classifies existing models based on criteria such as the sequence representation strategy, the underlying framework, or the target task. Our categorization aims to provide a structured overview of how these models are utilized across various vision tasks. We then present both quantitative and qualitative metrics to assess their performance and applicability. Finally, we highlight the current limitations of autoregressive models, such as computational complexity and mode collapse, and propose potential directions for future research. In summary, this survey makes several contributions:
@@
RQ-VAE (Lee et al., 2022a), MoVQ (Zheng et al., 2022), DQ (Huang et al., 2023b)
@@
DCT-Transformer (Nash et al., 2021), , fill=orange!10, text width=34em, align=left] [VQ-VAE (Van Den Oord et al., 2017), VQ-VAE-2 (Razavi et al., 2019),
@@
ARM (Ren et al., 2024b), AiM (Li et al., 2024b), , fill=orange!10, text width=34em, align=left] ] [Scale-wise Generation3.1.1.13), fill=orange!10 , text width=20em [VAR (Tian et al., 2024) , fill=orange!10, text width=34em] ] ] [Text-to-Image Synthesis3.1.2), fill=green!1 , text width=20em [Token-wise Generation6), fill=green!1, text width=20em [DALL·E (Ramesh et al., 2021), CogView (Ding et al., 2021),
@@
MVG (Ren et al., 2024a), fill=yellow!10, text width=34em] ] ] ] [Video Generation3.2), fill=pink!5 , text width=15em [Unconditional Video Generation3.2.1), fill=pink!5, text width=20em [MAGVIT (Yu et al., 2023a), MAGVIT-v2 (Yu et al., 2023b)
Open-MAGVIT2 (Luo et al., 2024), OmniTokenizer (Wang et al., 2024b) , fill=pink!5, text width=34em, align=left] [VPNs (Kalchbrenner et al., 2017), MoCoGAN (Tulyakov et al., 2018),
@@
ST-LLM (Liu et al., 2025), Pyramid Flow (Jin et al., 2024a), fill=pink!5, text width=34em, align=left] [MAGE (Hu et al., 2022), VideoPoet (Kondratyuk et al., 2024) , fill=pink!5, text width=34em] ] [Embodied AI3.2.3), fill=pink!5 ,text width=20em [IRIS (Micheli et al., 2023), GAIA-1 (Hu et al., 2023), iVideoGPT (Wu et al., 2024d),
@@
ImAM (Qian et al., 2024), Argus3D (Qian et al., 2024) , fill=violet!10, text width=34em ] ] [Scene Generation3.3.3), fill=violet!10, text width=20em [ SceneScript (Avetisyan et al., 2024) , fill=violet!10, text width=34em ] ] [3D Medical Generation3.3.4), fill=violet!10, text width=20em [SynthAnatomy (Tudosiu et al., 2022), BrainSynth (Tudosiu et al., 2024),
@@
Unalign (Corona-Figueroa et al., 2023), AutoSeq (Wang et al., 2024c), fill=violet!10, text width=34em, align=left ] ] ] [Multimodal Understanding and Generation3.4), fill=teal!7, text width=26em [Understanding Framework3.4.1), fill=teal!7, text width=20em [BEiT  (Bao et al., 2021), BEiT-v2 (Peng et al., 2022), LLaVA  (Liu et al., 2024d),
VL-BEiT (Bao et al., 2022), BEiT-v3 (Wang et al., 2022b)
AIM (El-Nouby et al., 2024), AIMV2 (Fini et al., 2024)
, fill=teal!7, text width=34em, align=left ] ] [
Unified Framework3.4.2), fill=teal!7, text width=20em [OFA (Wang et al., 2022a), CogView (Ding et al., 2021), M6 (Lin et al., 2021),
@@
Video-LaViT (Jin et al., 2024c), X-ViLA (Ye et al., 2024), fill=teal!7, text width=34em, align=left] [Chameleon (Team, 2024), Transfusion (Zhou et al., 2024)
@@ 2.1 Preliminary
The visual data is first transformed into a sequence of discrete elements. These elements may correspond to pixels, image patches, or latent codes derived from visual content, depending on the specific model architecture. This transformation allows the visual data to be framed as a sequential modeling problem, akin to text generation in natural language processing.
@@ 2.2 Generic Frameworks
In the preceding sections, we have provided a general overview of visual autoregressive models about the essential concept of sequence representation and sequence modeling. Then, we consider the representation strategy to classify the visual autoregressive models further, i.e. pixel-based models, token-based models, and scale-based models. Since representation strategy directly reflects how models handle the underlying structure and features of visual data. Different representation strategies influence the model’s granularity, efficiency, and applicability. Pixel-based models (Sec. 2.2.1) generate images pixel by pixel, capturing fine-grained spatial details. Token-based models (Sec. 2.2.2) represent images as a set of high-level discrete visual tokens, analogous to words in NLP. Scale-based models (Sec. 2.2.3) generate images through a multi-scale representation, allowing the model to handle visual content from different resolutions. We detail these three types of generic frameworks below.
2.2.0.1 Discussion about continuous methods
Our categorization of autoregressive generation methods is based on the representation strategy of the autoregressive elements. Although our discussion primarily focuses on discrete autoregressive methods, we note that continuous autoregressive approaches (Li et al., 2024c; Zhou et al., 2024), which operate directly in the embedding space without discretization, can be naturally accommodated within this framework. Continuous autoregressive models share many similarities with their discrete counterparts, with two notable differences: (a) they perform regression directly on continuous embeddings, eliminating the need for quantization; and (b) while discrete models typically minimize negative log-likelihood using a cross-entropy loss, continuous models require alternative loss formulations (e.g., diffusion loss (Li et al., 2024c)) on a case-by-case basis. Despite these differences, the underlying principles of autoregressive generation paradigms remain analogous, and our taxonomy is flexible enough to encompass both discrete and continuous methods.
@@ 2.2.2 Token-based Models
This quantization process enables a compact and discrete representation of the latent feature space, which is particularly advantageous for reducing the computational burden for high-dimensional image generation tasks. A seminal work on discrete representation is VQVAE (Van Den Oord et al., 2017), which introduced a two-stage paradigm that has become the de-facto standard for autoregressive vision generation. In this paradigm, an encoder-decoder architecture is initially trained to learn the discrete image representation. The encoder $E$ maps the input image $x$ to a latent space $z_{e}(x)$, where the continuous latent features are quantized into discrete codes using VQ. The decoder $D$ then reconstructs the image from these discrete codes. The training objective consists of a reconstruction loss and a commitment loss to ensure the latent codes effectively represent the input:
@@ 2.2.3 Scale-based Models
2.2.4 Analysis of Computational Costs
We compare the time complexity and efficiency of three autoregressive generation paradigms: next-pixel prediction, next-token prediction, and next-scale prediction. We consider a baseline “vanilla” architecture comprising a standard self-attention Transformer for autoregressive modeling, and an optional CNN-based tokenizer for decoding. Our analysis assumes greedy sampling without any additional efficiency techniques.
In next-pixel prediction, an $N\times N$ image is generated sequentially, with the Transformer’s self-attention incurring a per-step cost of $O_{T}(i^{2})$ for a sequence of length I. Consequently, generating all $N^{2}$ pixels requires:
$\sum_{i=1}^{N^{2}}i^{2}=\frac{1}{6}N^{2}(N^{2}+1)(2N^{2}+1)$ (6)
which is equivalent to $O_{T}(n^{6})$ basic computation. Next-token prediction reduces this cost by incorporating an image tokenizer with a compression ratio $k$, which shortens the latent sequence to $(N/k)^{2}$ tokens. The autoregressive cost is thereby reduced to $O_{T}(N^{6}/k^{6})$, and decoding the latent representation back to an $N\times N$ image via a CNN incurs a cost of $O_{c}(N^{2})$. The total cost is significantly lower than that of next-pixel prediction, particularly for high-resolution image synthesis. Next-scale prediction further enhances efficiency through a block-wise causal masking strategy and a reduced number of iterative steps, resulting in a time complexity of $O_{T}(N^{4}/k^{4})+O_{c}(N^{2})$ (Tian et al., 2024). Table 1 summarizes comparative computational complexities and efficiency ratings of these paradigms.
Require Tokenizer Compression Ratio Complexity Efficiency
Next-Pixel Prediction - $O_{T}(N^{6})$
Next-Token Prediction $k$ $O_{T}(N^{6}/k^{6})+O_{C}(N^{2})$
Next-Scale Prediction $k$ $O_{T}(N^{4}/k^{4})+O_{C}(N^{2})$
Table 1: Efficiency Comparison of Different Autoregressive Generation Paradigms. We consider the task of generating an $N\times N$ image using a standard self-attention Transformer and an optional CNN-based tokenizer.
@@ 2.3 Relation to Other Generative Models
Variational Autoencoders (VAEs) (Kingma, 2013) are a class of generative models that learn to map data into a continuous, lower-dimensional latent space and subsequently reconstruct it back to the original data space. This process is governed by optimizing a variational lower bound of the data likelihood. In contrast, autoregressive models directly capture the full data distribution by predicting each element sequentially, optimizing negative log-likelihood(NLL) as the training objective. While VAEs provide efficient sampling, their reliance on variational approximations often results in less sharp outputs (Child, 2021) and may suffer from posterior collapse (Zhao et al., 2019; Chen et al., 2017) when equipped with a powerful decoder. Autoregressive models, on the other hand, generate high-quality samples by modeling data dependencies at the original dimensionality, but they are slower in inference due to their sequential generation process. To address these limitations, hybrid approaches (Gulrajani et al., 2017; Van Den Oord et al., 2017; Chen et al., 2017) have been proposed that integrate the autoregressive process into VAEs. One prominent example is VQ-VAE (Van Den Oord et al., 2017), which utilizes a VAE framework to learn discrete latent spaces and autoregressive models to refine generation, effectively leveraging the strengths of both models for efficient and high-quality image synthesis.
@@
GANs (Goodfellow et al., 2014) are known for generating high-quality images with fast inference, enabled by their one-shot generation process. GANs demonstrate particular strength in domains with structured data, such as facial synthesis, where their low-dimensional latent space can effectively manipulate visual attributes (Xia et al., 2022). However, the adversarial training objective that GANs rely on often results in training instability and mode collapse (Liu et al., 2020), which requires a delicate balance between the generator and discriminator. Additionally, the adversarial training paradigm that underpins GANs can hinder their scalability to more diverse datasets and larger model sizes. In contrast, autoregressive models employ likelihood-based training, which ensures a stable training process. Despite their relatively slow sampling speeds, autoregressive models exhibit favorable scaling laws, where model performance consistently improves with larger datasets and increased model sizes (Kaplan et al., 2020). These properties make autoregressive models particularly well-suited for building universal generative models and can be adapted to a wide range of downstream tasks through zero-shot generalization of supervised fine-tuning (SFT), enabling flexibility and robustness across different applications.
@@
Diffusion models (Ho et al., 2020) have emerged as the state-of-the-art generation paradigm for a wide range of vision generation tasks. As recent challengers, autoregressive models share several characteristics with diffusion models. Both methods can generate diverse, high-quality samples, yet both also suffer from inefficient inference due to their iterative or sequential generation processes. Additionally, both approaches are likelihood-based and optimize objectives such as Negative Log-Likelihood (NLL) or Evidence Lower Bound (ELBO), making them relatively easy to train. Recent advances (Esser et al., 2021b; Rombach et al., 2022) in both fields have focused on compressing visual content into latent spaces to improve the efficiency of high-resolution generation. Despite these similarities, there are fundamental differences in their generative paradigms. Diffusion models employ a predefined forward process that gradually corrupts data and an iterative denoising process to recover samples. This iterative process retains spatial locality at each step, which benefits visual coherence. However, the necessity to corrupt the training data with Gaussian noise may potentially limit their scalability for understanding tasks. Additionally, while diffusion models operate primarily in continuous spaces, their discrete variants still lag behind in performance, posing challenges for multimodal applications like text generation. Autoregressive models, by contrast, inherently introduce unidirectional biases and discretization, which might not be ideal for visual tasks, partially explaining their lag behind diffusion models in recent benchmarks. However, autoregressive models offer flexibility in handling diverse modalities and combining generation with understanding. This adaptability aligns with the emerging trend of integrating autoregressive models with Large Language Models (LLMs) to create a unified framework for multimodal input-output tasks, bridging both generation and understanding capabilities. These distinctions and complementarities between the two approaches suggest potential avenues for future research. Recent efforts have aimed to combine the strengths of both approaches, such as integrating bidirectional contexts (Esser et al., 2021a; Zhou et al., 2024; Xie et al., 2024b), autoregressive models with continuous representations (Li et al., 2024c; Zhou et al., 2024), discrete diffusion models (Gu et al., 2022), and using autoregressive models to enhance the understanding capabilities of diffusion models (Gu et al., 2024).
@@
2.4 Comparison between Autoregressive Methods and Non-Autoregressive Methods
In this subsection, we discuss the unique advantages of autoregressive models for vision in comparison to non-autoregressive models, such as the popular diffusion-based (Ho et al., 2020) and rectified-flow based models (Liu et al., 2022). We summarize these advantages in three key points:
1. Scaling Laws: The success of the next-token prediction paradigm in NLP (Achiam et al., 2023; Lyu et al., 2023; Dubey et al., 2024) can be largely attributed to well-established scaling laws (Henighan et al., 2020; Kaplan et al., 2020). Although the scaling properties of methods like diffusion and GANs remain relatively underexplored, autoregressive visual models offer the potential to transfer successful scaling experiences from NLP. This could enable efficient scaling in visual generation frameworks, particularly in terms of model size and performance. Some work like VAR (Tian et al., 2024), LlamaGen (Sun et al., 2024a) and FLUID (Fan et al., 2024) exemplify how autoregressive methods benefit from these scaling laws, highlighting their potential for advancing visual generation.
2: Deployment Efficiency: Autoregressive generative models can leverage the existing deployment technologies designed for language models (Wan et al., 2023). Frameworks such as VLLM provide acceleration for autoregressive models, significantly improving generation efficiency. This allows autoregressive visual models to benefit from infrastructure that has already been optimized for language tasks, facilitating seamless deployment in real-world applications.
3. Bridging Language and Vision: Autoregressive models offer a promising pathway toward unifying multimodal understanding and generation. By aligning closely with the structure of large language models, autoregressive approaches may provide a natural bridge between vision and language tasks. This synergy opens up the potential for more integrated and versatile models that can handle both vision and language processing in a unified manner, contributing to a deeper understanding of multimodal tasks (Team, 2024; Zhou et al., 2024; Xie et al., 2024b; Wu et al., 2024a).
In addition to these primary advantages, autoregressive models share some secondary benefits with certain non-autoregressive models. For example, autoregressive training demonstrates inherent stability. By directly optimizing the negative log-likelihood (NLL) through the minimization of the cross-entropy loss function, as depicted in Eq. 2, autoregressive models eliminate the need for adversarial training, as is the case with GANs (Goodfellow et al., 2014), or for the complex design of invertible networks like normalizing flows (Dinh et al., 2014; Rezende & Mohamed, 2015). Furthermore, autoregressive models often demonstrates well generation diversity, less worrying about the model collapse curse that can affect non-autoregressive models.
Despite these advantages, autoregressive models do have certain limitations. One significant drawback, when compared to diffusion models, is their difficulty in generating ultra-high-resolution images or videos. The higher-order time complexity of autoregressive generation makes this task more challenging. As discussed in Table 1, autoregressive models using next-token prediction paradigm have a time complexity of $O(N^{6}/k^{6})$ when generating an $N\times N$ image, whereas a Diffusion Transformer (DiT) model (Peebles & Xie, 2023) in a similar setting typically requires only $O(TN^{4})$ time complexity (T for the denoising steps). Additionally, while autoregressive generation has shown promising progress in prompt-following when generating images from given prompts (Fan et al., 2024; Wang et al., 2024d), as demonstrated by benchmarks like GenEval (Ghosh et al., 2024), the visual quality of single images generated by autoregressive models still lags behind state-of-the-art diffusion or flow-based models (Esser et al., 2024; Labs, 2024), and significant research is still needed to bridge this gap. We provide a more detailed discussion of potential future directions for this work in Section 5.